# Bogomolov property and Galois representations

### v1

@@ Abstract.
In 2013 P. Habegger proved the Bogomolov property for the field generated over ${\mathbb{Q}}$ by the torsion points of a rational elliptic curve. We explore the possibility of applying the same strategy of proof to the case of field extensions fixed by the kernel of some modular Galois representations.
${}^{(1)}$ipartimento di Matematica
@@
Via Carlo Alberto, 10 - 10123 Torino
${}^{(2)}$ipartimento di Informatica
@@
Corso Svizzera, 185 - 10149 Torino
@@ 1. Introduction
Let $\alpha$ be an algebraic number of degree $d$. We denote by $h(\alpha)$ the absolute and logarithmic Weil height. Assuming that $\alpha$ is not a root of unity, D. H. Lehmer suggested that $dh(\alpha)$ can be bounded below uniformly in $\alpha$. Known as “Lehmer’s problem”, this question is still open. Nevertheless, in some special cases, not only Lehmer’s conjecture is true, but it can also be sharpened. This different point of view was considered for the first time in [5], where the authors introduced the notion of fields with Bogomolov property, denoted Property (B), where the height is uniformly bounded below outside torsion points. By Northcott’s Theorem, number fields are trivial examples of fields with (B). The compositum ${\mathbb{Q}}^{\rm tr}$ of all totally real fields is a less trivial example (apply [22, Corollary $1^{\prime}$] to the linear polynomial $P(z)=z-\alpha$ to get the optimal lower bound). In [5] the authors also show that fields with bounded local degrees at some finite place has Property (B). Note that by a result of [8] a Galois extension has uniformly bounded local degrees at every finite place if and only if its Galois group has finite exponent. Thus (infinite) Galois extensions with Galois group of bounded exponent have Property (B).
@@
There are still other several interesting examples of subfields of the algebraic numbers with Property (B). Zannier conjectured that the maximal abelian extension ${\mathbb{Q}}^{\rm ab}$ of the rational field also satisfies property (B), and this was proved in [2]. By a special case of [3, Theorem 1.1], the same is true for $K^{\rm ab}$, for any number field $K$. Moreover we can combine this result with the quoted result on Galois extensions of bounded exponent. Let $K$ be a number field and let $L/K$ be an infinite Galois extension with Galois group $G$ with $G/Z(G)$ of finite exponent. Then $L$ has Property (B) ([1, Corollary 1.7]).
@@
Another class of fields with Property (B) has been exhibited in [15]. Let $E$ be an elliptic curve defined over ${\mathbb{Q}}$. Then the field ${\mathbb{Q}}(E_{tors})$ obtained by adjoining all torsion points of $E$ has the Bogomolov property. If $E$ has complex multiplication then ${\mathbb{Q}}(E_{tors})$ is an abelian extension of a quadratic number field and thus it has the property (B) by the quoted result of [3]. However for a non-CM elliptic curve there are no number fields K such that ${\mathbb{Q}}(E_{tors})$ is contained in the maximal abelian extension of K.
@@
Note that by Kronecker-Weber theorem ${\mathbb{Q}}^{\rm ab}$ is the field obtained by adjoining all torsion points of the multiplicative group. A conjecture due to S. David (see [13, Conjecture, p.114]) suggests that ${\mathbb{Q}}(A_{tors})$ has Property (B), when $A$ is an abelian variety defined on a number field, according as it is already known for the algebraic groups ${\mathbb{G}}_{\rm m}$ and $E=$ elliptic curve. Note that for a CM abelian variety $A$ the field ${\mathbb{Q}}(A_{tors})$ is still an abelian extension of a (CM) number field, and thus property (B) holds, again by the quoted result of [3].
@@
Habegger’s method provides a lower bound $h(\alpha)\geq c(p)$ with $c(p)>0$ whenever we have a prime $p\geq 5$ such that $E$ has supersingular reduction at $p$ (assumption (P1)) and the Galois representation $\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}})\to{\rm Aut}E[p]$ is surjective (assumption (P2)). See [13, Theorem 1.3, p.114] for an explicit result. By a result of Elkies there exist infinitely many primes satisfying (P1); and by a result of Serre (P2) holds outside a finite set of primes. Thus we can find a prime $p$ satisfying both the above conditions. Let now A be an abelian variety. Even under similar assumptions for a prime $p$, an analogue of the result of Habegger is not known.
@@
Let $\Omega$ be a topological group such that points are closed, $F$ be a subfield of $\overline{\mathbb{Q}}$ and $G_{F}=\mathrm{Gal}(\overline{\mathbb{Q}}/F)$. We say that a continuous homomorphism
$\rho:G_{F}\longrightarrow\Omega,$
has Property (B) if the field $\mathbb{Q}(\rho):=\overline{\mathbb{Q}}^{\ker(\rho)}$ does.
@@
In this vein, [1, Corollary 1.7] has the following consequence. Let $K$ be a number field, $\rho:G_{K}\to GL_{n}(F)$ be an absolutely irreducible representation, and $L=\mathbb{Q}(\rho)$. Schur’s Lemma asserts that every intertwining automorphism of $\rho$ is a scalar operator. In particular the image of the center of $G=\mathrm{Gal}(L/K)$ is contained in the subgroup of scalar matrices of $\mathrm{GL}_{2}(F)$, so that the reduction map $\tilde{\rho}:G\to\mathrm{PGL}_{n}(F)$ factors through $G/Z(G)$. Thus, if the image of $\tilde{\rho}$ has finite exponent, the same is true for $G/Z(G)$ and [1, Corollary 1.7] implies that $\rho$ has property (B).
@@
$\rho_{E}:G_{\mathbb{Q}}\longrightarrow GL_{2}(\widehat{\mathbb{Z}})$
@@
$\rho_{E,\ell}:\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}})\longrightarrow Aut(T_{\ell}(E))=GL_{2}(\mathbb{Z}_{\ell})$
where $E$ is an elliptic curve defined over ${\mathbb{Q}}$ and $T_{\ell}(E)$ is the corresponding Tate module. Indeed, by Galois correspondence ${\mathbb{Q}}(E[\ell^{n}])$ is the subfield of $\overline{\mathbb{Q}}$ fixed by the kernel of the action of $G_{\mathbb{Q}}$ on the $\ell^{n}$-torsion points of $E$; by passing to the projective limit we see that ${\mathbb{Q}}(E[\ell^{\infty}])=\mathbb{Q}(\rho_{E,\ell})$. Since ${\mathbb{Q}}(E_{tors})$ is the compositum of ${\mathbb{Q}}(E[\ell^{\infty}])$, for every prime $\ell$, then it coincides with the field fixed by $\bigcap_{\ell}\ker(\rho_{E,\ell})=\ker(\rho_{E})$.
By the Shimura-Taniyama-Weil conjecture, proved by Wiles, all elliptic curves defined over ${\mathbb{Q}}$ are associated to a cuspidal eigenform of weight $2$ on $\Gamma_{0}(N)$, whose eigenvalues for the Hecke operators are rational numbers. A question that naturally arises starting from Habegger’s theorem is whether and to what extent it can be generalized to other Galois representations associated with modular forms. Let $k\geq 2$ be a positive integer and $f\in S_{k}(\Gamma_{0}(N))$ be a normalized eigenform for the Hecke algebra. Let $K_{f}$ be the number field generated by the Hecke eigenvalues of $f$, and let $\mathcal{O}_{f}$ be its ring of integers. A construction due to Shimura and Deligne attaches to $f$ a family of compatible Galois representations
$\rho_{f,v}:G_{\mathbb{Q}}\longrightarrow GL_{2}(\mathcal{O}_{f,v})$
@@
$\rho_{f,p}:G_{\mathbb{Q}}\longrightarrow\prod_{v|p}GL_{2}(\mathcal{O}_{f,v}).$
@@
$\rho_{f}=\prod_{v}\rho_{f,v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\hat{\mathcal{O}}_{f}),$
where $\hat{\mathcal{O}}_{f}$ is the profinite completion of $\mathcal{O}_{f}$.
@@
When $k=2$, Shimura [25] associated to $f$ an abelian variety $A_{f}$ defined over ${\mathbb{Q}}$, of dimension $[K_{f}:{\mathbb{Q}}]$ equipped with an action of $\mathcal{O}_{f}$; such abelian varieties have been called of type $\mathrm{GL}_{2}$ in [21] and a generalization of Wiles’ theorem [17, Corollary 10.2] establishes that all of them arise from modular forms.
@@
${\mathbb{Q}}\left(\rho\right)={\mathbb{Q}}\left(\prod_{p}\rho_{p}\right)=\coprod_{p}{\mathbb{Q}}\left(A[p^{\infty}]\right)={\mathbb{Q}}\left(\bigcup_{p}A[p^{\infty}]\right)={\mathbb{Q}}\left(\sum_{p}A[p^{\infty}]\right),$
@@
Let $f=\sum_{n\geq 1}a_{n}q^{n}\in S_{2}(\Gamma_{0}(N))$ be a normalized eigenform. Assume that there exists a rational prime $p\geq 5$ such that
  • $(P0)$
    $p\nmid N$
  • $(P1)$
    $a_{p}=0$
  • $(P2)$
    $\widehat{G}(\mathcal{O}_{f}/p\mathcal{O}_{f})=\{A\in\mathrm{GL}_{2}(\mathcal{O}_{f}/p\mathcal{O}_{f})\ |\ \det(A)\in\mathbb{F}_{p}^{\times}\}.$
Let $F/\mathbb{Q}$ be a Galois extension unramified at $p$ with property (B). Then $\rho_{f,p}|_{G_{F}}$ has property (B).
As a consequence (see Corollary 3.5 for a proof), we obtain the following direct generalization of Habegger’s result:
@@
Let $f=\sum_{n\geq 1}a_{n}q^{n}\in S_{2}(\Gamma_{0}(N))$ be a normalized eigenform. Assume that there exists a prime $p\geq 5$ satisfying properties (P0), (P1), (P2) in Theorem 1.3. Then $\rho_{f}$ has property (B).
@@
Recall that a modular form $f$ is supersingular at a prime $\mathfrak{p}$ of $\mathcal{O}_{f}$ above $p$ if $a_{p}\in\mathfrak{p}$. When $a_{p}\in\mathbb{Z}$ (as in the case $K_{f}={\mathbb{Q}}$) and $p\geq 5$, this implies $a_{p}=0$, by the Deligne bound: $|a_{p}|\leq 2\sqrt{p}$. Therefore in the case of elliptic curves, Habegger’s assumption (P1) ($p\geq 5$ supersingular) is equivalent to $a_{p}=0$. Note that when $K_{f}\not={\mathbb{Q}}$ the condition $a_{p}=0$, required in our proof, is in general stronger than supersingularity.
@@
Our condition (P2) generalizes Habegger’s assumption on the surjectivity of the Galois representation $G_{\mathbb{Q}}\to{\rm Aut}E[p]$. We also remark that property (P2) has been extensively investigated in the literature (see [20, 21, 4, 14, 18]).
@@
Here we only limit ourselves to observe that the results in [20] imply (P2) when $k=2$, $N$ is a prime, $p\nmid 6(N-1)$, $K_{f}$ is unramified at $p$ and $\mathcal{O}_{f}$ is generated over $\mathbb{Z}$ by the coefficients $a_{n}$.
@@
Some explicit examples of such modular forms satisfying also (P1) (and (P0)) are the following:
  • $N=73$, $p=59$, $K_{f}=\mathbb{Q}(\sqrt{13})$ (form 73.2.a.c in [26]).
  • $N=167$, $p=11$, $K_{f}=\mathbb{Q}(\sqrt{5})$ (form 167.2.a.a in [26]).
  • $N=383$, $p=13$, $K_{f}=\mathbb{Q}(\sqrt{5})$ (form 383.2.a.a in [26].
  • $N=151$, $p=41$, $[K_{f}:\mathbb{Q}]=3$ (form 151.2.a.a in [26]).
We remark the generalization of Theorems 1.3 and 1.4 to the case $k>2$ seems to hold true as well (under the additional condition $p\geq 2k-1$), although the proof presents some further technical complications. It will be dealt with in another paper.
2. Diophantine results.
In this section we propose two general statements on lower bounds for the height. The proofs relies on the techniques developed in [2], [3], [1] and [15].

### v3

@@ Abstract.
In 2013 P. Habegger proved the Bogomolov property for the field generated over ${\mathbb{Q}}$ by the torsion points of a rational elliptic curve. We explore the possibility of applying the same strategy of proof to the case of field extensions cut out by Galois representations arising from more general modular forms.
(1)Dipartimento di Matematica
@@
Via Carlo Alberto, 10 - 10123 Torino, Italy
(2)Dipartimento di Informatica
@@
Corso Svizzera, 185 - 10149 Torino, Italy
2020 Mathematical Subject Classification: 11F03, 11F80, 11G50
Keywords. Weil height, Bogomolov property, modular forms, Galois representations.
@@ 1. Introduction
Let $\alpha$ be an algebraic number of degree $d$. We denote by $h(\alpha)$ the absolute and logarithmic Weil height. Assuming that $\alpha$ is not a root of unity, D. H. Lehmer suggested that $dh(\alpha)$ can be bounded below uniformly in $\alpha$. Known as “Lehmer’s problem”, this question is still open. Nevertheless, in some special cases, not only Lehmer’s conjecture is true, but it can also be sharpened. This different point of view was considered for the first time in [5], where the authors introduced the notion of fields with Bogomolov property, (property (B) for short), where the height is uniformly bounded below outside torsion points. By Northcott’s Theorem, number fields are trivial examples of fields with (B). The compositum ${\mathbb{Q}}^{\rm tr}$ of all totally real fields is a less trivial example (apply [24, Corollary $1^{\prime}$] to the linear polynomial $P(z)=z-\alpha$ to get the optimal lower bound). In [5] the authors also show that fields with bounded local degrees at some finite place has Property (B). Note that by a result of [9] a Galois extension has uniformly bounded local degrees at every finite place if and only if its Galois group has finite exponent. Thus (infinite) Galois extensions with Galois group of bounded exponent have Property (B).
@@
There are still other several interesting examples of subfields of the algebraic numbers with Property (B). Zannier conjectured that the maximal abelian extension ${\mathbb{Q}}^{\rm ab}$ of the rational field also satisfies property (B), and this was proved in [2]. By a special case of [3, Theorem 1.1], the same is true for ${\mathbb{K}}^{\rm ab}$, for any number field ${\mathbb{K}}$. Moreover we can combine this result with the quoted result on Galois extensions of bounded exponent. Let ${\mathbb{K}}$ be a number field and let ${\mathbb{L}}/{\mathbb{K}}$ be an infinite Galois extension with Galois group $G$ with $G/Z(G)$ of finite exponent. Then ${\mathbb{L}}$ has Property (B) ([1, Corollary 1.7]).
@@
Another class of fields with Property (B) has been exhibited in [15]. Let ${\mathcal{E}}$ be an elliptic curve defined over ${\mathbb{Q}}$. Then the field ${\mathbb{Q}}({\mathcal{E}}_{tors})$ obtained by adjoining all torsion points of ${\mathcal{E}}$ has the Bogomolov property. If ${\mathcal{E}}$ has complex multiplication then ${\mathbb{Q}}({\mathcal{E}}_{tors})$ is an abelian extension of a quadratic number field and thus it has the property (B) by the quoted result of [3]. However for a non-CM elliptic curve there are no number fields ${\mathbb{K}}$ such that ${\mathbb{Q}}({\mathcal{E}}_{tors})$ is contained in the maximal abelian extension of ${\mathbb{K}}$.
@@
Note that by Kronecker-Weber theorem ${\mathbb{Q}}^{\rm ab}$ is the field obtained by adjoining all torsion points of the multiplicative group. A conjecture due to S. David (see [13, Conjecture, p.114]) suggests that ${\mathbb{Q}}(A_{tors})$ has Property (B), when $A$ is an abelian variety defined on a number field, according as it is already known for the algebraic groups ${\mathbb{G}}_{\rm m}$ and ${\mathcal{E}}=$ elliptic curve. Note that for a CM abelian variety $A$ the field ${\mathbb{Q}}(A_{tors})$ is still an abelian extension of a (CM) number field, and thus property (B) holds, again by the quoted result of [3].
@@
Habegger’s method provides a lower bound $h(\alpha)\geq c(p)$ with $c(p)>0$ whenever we have a prime $p\geq 5$ such that ${\mathcal{E}}$ has supersingular reduction at $p$ (assumption $(P1)$) and the Galois representation $\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}})\to{\rm Aut}({\mathcal{E}}[p])$ is surjective (assumption $(P2)$). See [13, Theorem 1.3, p.114] for an explicit result. By a result of Elkies there exist infinitely many primes satisfying $(P1)$; and by a result of Serre $(P2)$ holds outside a finite set of primes. Thus we can find a prime $p$ satisfying both the above conditions. Let now A be an abelian variety. Even under similar assumptions for a prime $p$, an analogue of the result of Habegger is not known.
@@
Let $\Omega$ be a separated topological group, ${\mathbb{F}}$ be a subfield of $\overline{\mathbb{Q}}$ and $G_{\mathbb{F}}=\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{F}})$. We say that a continuous homomorphism
$\rho:G_{\mathbb{F}}\longrightarrow\Omega,$
has property (B) if $\mathbb{F}(\rho):=\overline{\mathbb{Q}}^{\ker(\rho)}$ does.
@@
In this vein, [1, Corollary 1.7] has the following consequence. Let ${\mathbb{K}}$ be a number field, $\rho:G_{\mathbb{K}}\to\mathrm{GL}_{n}({\mathbb{F}})$ be an absolutely irreducible representation, and ${\mathbb{L}}=\mathbb{K}(\rho)$. Schur’s Lemma asserts that every intertwining automorphism of $\rho$ is a scalar operator. In particular the image of the center of $G=\mathrm{Gal}({\mathbb{L}}/{\mathbb{K}})$ is contained in the subgroup of scalar matrices of $\mathrm{GL}_{n}({\mathbb{F}})$, so that the reduction map $\tilde{\rho}:G\to\mathrm{PGL}_{n}({\mathbb{F}})$ factors through $G/Z(G)$. Thus, if the image of $\tilde{\rho}$ has finite exponent, the same is true for $G/Z(G)$ and [1, Corollary 1.7] implies that $\rho$ has property (B).
@@
$\rho_{\mathcal{E}}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\widehat{\mathbb{Z}})$
@@
$\rho_{{\mathcal{E}},\ell}:\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}})\longrightarrow Aut(T_{\ell}({\mathcal{E}}))=\mathrm{GL}_{2}(\mathbb{Z}_{\ell})$
where ${\mathcal{E}}$ is an elliptic curve defined over ${\mathbb{Q}}$ and $T_{\ell}({\mathcal{E}})$ is the corresponding Tate module. Indeed, by Galois correspondence ${\mathbb{Q}}({\mathcal{E}}[\ell^{n}])$ is the subfield of $\overline{\mathbb{Q}}$ fixed by the kernel of the action of $G_{\mathbb{Q}}$ on the $\ell^{n}$-torsion points of ${\mathcal{E}}$; by passing to the projective limit we see that ${\mathbb{Q}}({\mathcal{E}}[\ell^{\infty}])=\mathbb{Q}(\rho_{{\mathcal{E}},\ell})$. Since ${\mathbb{Q}}({\mathcal{E}}_{tors})$ is the compositum of ${\mathbb{Q}}({\mathcal{E}}[\ell^{\infty}])$, for every prime $\ell$, then it coincides with the field fixed by $\bigcap_{\ell}\ker(\rho_{{\mathcal{E}},\ell})=\ker(\rho_{{\mathcal{E}}})$.
By the Shimura-Taniyama-Weil conjecture, proved by Wiles, all elliptic curves defined over ${\mathbb{Q}}$ are associated to a cuspidal eigenform of weight $2$ on $\Gamma_{0}(N)$, whose eigenvalues for the Hecke operators are rational numbers. A question that naturally arises starting from Habegger’s theorem is whether and to what extent it can be generalized to other Galois representations associated with modular forms. Let $k\geq 2$ be a positive integer and $f\in S_{k}(\Gamma_{0}(N))$ be a normalized eigenform for the Hecke algebra. Let ${\mathbb{K}}_{f}$ be the number field generated by the Hecke eigenvalues of $f$, and let $\mathcal{O}_{f}$ be its ring of integers. A construction due to Shimura and Deligne attaches to $f$ a family of compatible Galois representations
$\rho_{f,v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\mathcal{O}_{f,v})$
@@
$\rho_{f,p}:G_{\mathbb{Q}}\longrightarrow\prod_{v|p}\mathrm{GL}_{2}(\mathcal{O}_{f,v}).$
@@
$\rho_{f}=\prod_{v}\rho_{f,v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\widehat{\mathcal{O}}_{f}),$
where $\widehat{\mathcal{O}}_{f}$ is the profinite completion of $\mathcal{O}_{f}$.
@@
If $f$ is a CM modular form, then $\mathbb{Q}(\rho_{f})$ is an abelian extension of a quadratic field, (see for instance [20, Proposition 4.4]). Thus $\rho_{f}$ has property (B) by the special case $D=1$ of [3, Theorem 1.1].
When $k=2$, Shimura [28] associated to $f$ an abelian variety $A$ defined over ${\mathbb{Q}}$, of dimension $[{\mathbb{K}}_{f}:{\mathbb{Q}}]$ equipped with an action of $\mathcal{O}_{f}$; such abelian varieties have been called of type $\mathrm{GL}_{2}$ in [23] and a generalization of Wiles’ theorem [17, Corollary 10.2] establishes that all of them arise from modular forms.
@@
${\mathbb{Q}}\left(\rho_{f}\right)={\mathbb{Q}}\left(\prod_{p}\rho_{f,p}\right)=\coprod_{p}{\mathbb{Q}}\left(A[p^{\infty}]\right)={\mathbb{Q}}\left(\bigcup_{p}A[p^{\infty}]\right)={\mathbb{Q}}\left(\sum_{p}A[p^{\infty}]\right),$
@@
Assume that $\sum_{n\geq 1}a_{n}q^{n}$ is the $q$-expansion of $f$. Let $p$ be a rational prime. Having fixed $f$, we consider the following assumption on $f$ and $p$:
Assumption 1.3.
  • $(P0)$
    $p\nmid N$;
  • $(P1)$
    $a_{p}=0$;
  • $(P2)$
    $\widehat{G}(\mathcal{O}_{f}/p\mathcal{O}_{f})=\{\gamma\in\mathrm{GL}_{2}(\mathcal{O}_{f}/p\mathcal{O}_{f})\ |\ \det\gamma\in(\mathbb{F}_{p}^{\times})^{k-1}\}.$
  • $(P3)$
     [^2] Note that all these conditions are satisfied if $p\geq 2k-1$. $p\geq 5$, $p\nmid k-1$ and $\frac{p+1}{2}\nmid k-1$.
@@
Let $f=\sum_{n\geq 1}a_{n}q^{n}\in S_{k}(\Gamma_{0}(N))$ be a normalized eigenform. Assume that there exists a rational prime $p$ which satisfies Assumption 1.3. Let ${\mathbb{F}}/\mathbb{Q}$ be a Galois extension unramified at $p$ with property (B). Then $\rho_{f,p}|_{G_{\mathbb{F}}}$ has property (B).
As a consequence (see section 6 for the proof), we obtain the following direct generalization of Habegger’s result:
@@
Let $f=\sum_{n\geq 1}a_{n}q^{n}\in S_{k}(\Gamma_{0}(N))$ be a normalized eigenform. Assume that there exists a rational prime $p$ satisfying Assumption 1.3. Then $\rho_{f}$ has property (B).
@@
Recall that a modular form $f$ is supersingular at a prime $\mathfrak{p}$ of $\mathcal{O}_{f}$ above $p$ if $a_{p}\in\mathfrak{p}$. When $a_{p}\in\mathbb{Z}$ (as in the case ${\mathbb{K}}_{f}={\mathbb{Q}}$) and $p\geq 5$, this implies $a_{p}=0$, by the Deligne bound: $|a_{p}|\leq 2\sqrt{p}$. Therefore in the case of elliptic curves, Habegger’s assumption $p\geq 5$ supersingular is equivalent to $a_{p}=0$. Note that when ${\mathbb{K}}_{f}\not={\mathbb{Q}}$ the condition $a_{p}=0$, required in our proof, is in general stronger than supersingularity.
@@
Our condition $(P2)$ generalizes Habegger’s assumption on the surjectivity of the Galois representation $G_{\mathbb{Q}}\to{\rm Aut}{\mathcal{E}}[p]$. It has been extensively investigated in the literature (see [22, 23, 4, 14, 18]); we summarize here the most relevant results for the purposes of this work.
In the case of modular forms associated to a non CM elliptic curve, a theorem of Serre [27] establishes property $(P2)$ for every $p$ except finitely many. This result, together with the aforementioned Elkies’theorem, ensures the non conditional character of Habegger’s theorem.
For a modular form such that $\mathbb{K}_{f}\not=\mathbb{Q}$ the situation concerning the image of $\rho_{f,p}$ is complicated by the possible of presence of inner twists; i.e. the elements of the subgroup $\Gamma\subseteq Aut(\mathbb{K}_{f})$ consisting of those automorphisms $\sigma$ for which $f^{\sigma}$ is equal to the twist of $f$ by some Dirichlet character. Let $\mathbb{K}_{1}$ be the subfield of $\mathbb{K}$ fixed by $\Gamma$. It is proved in [22, Theorem 3.1] that there is an abelian number field $E$ such that for $p$ large enough relatively to $f$, the restriction $\rho_{f,p}|_{G_{E}}$ has image $\widehat{G}(\mathcal{O}_{\mathbb{K}_{1}}\otimes\mathbb{Z}_{p})$.
Therefore, all we can say is that condition $(P2)$ holds for $p$ large enough when $f$ has no inner twists. In the case of trivial nebentypus (the one we are dealing with) it is easy to see that inner twists must arise from a quadratic Dirichlet character; and Ribet proved that they do not appear when $N$ is square free [21, 3.9].
@@
In particular, the results in [22] imply $(P2)$ when $k=2$, $N$ is a prime, $p\nmid 6(N-1)$, ${\mathbb{K}}_{f}$ is unramified at $p$ and $\mathcal{O}_{f}$ is generated over $\mathbb{Z}$ by the coefficients $a_{n}$.
@@
Some explicit examples with $k=2$ of such modular forms satisfying also $(P0)$ and $(P1)$ are the following:
  1. -
    $N=73$, $p=59$, ${\mathbb{K}}_{f}=\mathbb{Q}(\sqrt{13})$ (form 73.2.a.c in [29]).
  2. -
    $N=167$, $p=11$, ${\mathbb{K}}_{f}=\mathbb{Q}(\sqrt{5})$ (form 167.2.a.a in [29]).
  3. -
    $N=383$, $p=13$, ${\mathbb{K}}_{f}=\mathbb{Q}(\sqrt{5})$ (form 383.2.a.a in [29].
  4. -
    $N=151$, $p=41$, $[{\mathbb{K}}_{f}:\mathbb{Q}]=3$ (form 151.2.a.a in [29]).
Some other examples in weight $k>2$ with $a_{p}=0$, can be produced when $\mathbb{K}=\mathbb{Q}$. In this case in fact conditions (i),(ii), (iii) in [22, page 192] are automatically fulfilled, and if $a_{p}=0$, then condition (iv) is also satisfied, because the residual representation $\bar{\rho}_{v}$ is irreducible, for every $v|p$. There are several examples of this type with $k=4$, for example
  1. -
    $N=186$, $p=11$, (form 186.4.a.a in [29]).
  2. -
    $N=210$, $p=11$ or $p=23$, (form 210.4.a.e in [29]).
  3. -
    $N=1265$, $p=53$, (form 1265.4.a.c in [29]).
Some other examples with $k\geq 6$ are:
  1. -
    $N=390$, $p=7$, $k=6$ ((form 390.6.a.c in [29]).
  2. -
    $N=66$, $p=5$, $k=8$ (form 66.8.a.a in [29]).
We were not able to find examples with weight $>8$.
Indeed, the simultaneous occurrence of conditions $(P0)$ and $(P1)$ for a non CM modular form is expected to be a fairly rare phenomenon. In this direction, Calegari and Sardari [7] proved that, fixed a pair of coprime integers $p,N$, there is only a finite number of eigenforms for $\Gamma_{1}(N)$ having $a_{p}=0$.
Plan of the paper.
In Section 2 we introduce most of the notations we need and we make some preliminary remarks.